OpenStudy (anonymous):

solve for x to the nearest hundredth: 6^x=45

OpenStudy (solomonzelman):

6^x=45 hit both sides with log log(6^x)=log45 knowing that log(a^b)=b log a, xlog6=log45 x=log45/log6 can you do it now?